OpenStudy (anonymous):

Why can a subtraction problem be written as an addition problem without changing the outcome

OpenStudy (welshfella):

an example would be a - b = a + (-b)

OpenStudy (baru):

for example if your question was 35-10= (?) you can rephrase this question as 10 + (?) =35